# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

### v2

@@ Abstract
Retrieval-Augmented-Generation and Gener-ation-Augmented-Generation have been proposed to enhance the knowledge required for question answering over Large Language Models (LLMs). However, the former depends on external resources, and both require incorporating the explicit documents into the context, which results in longer contexts that lead to more resource consumption. Recent works indicate that LLMs have modeled rich knowledge, albeit not effectively triggered or activated. Inspired by this, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG), which simulates the human capacity to compensate for knowledge deficits while answering questions solely through imagination, without relying on external resources. Guided by IAG, we propose an imagine richer context method for question answering (IMcQA), which obtains richer context through the following two modules: explicit imagination by generating a short dummy document with long context compress and implicit imagination with HyperNetwork for generating adapter weights. Experimental results on three datasets demonstrate that IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in both in-distribution performance and out-of-distribution generalizations 111 Our code will be available at https://github.com/Xnhyacinth/IAG.
@@ 1 Introduction
Knowledge-intensive tasks like question answering (QA) necessitate access to extensive world and domain knowledge Berant et al. (2013); Joshi et al. (2017); Kwiatkowski et al. (2019). Recently, Large Language Models (LLMs) have displayed notable competencies in almost every task and industry within the “pre-train, prompt, and predict” paradigm Liu et al. (2023b). However, LLMs lack the sufficient capability to independently handle knowledge-intensive tasks Yu et al. (2023) and usually generate hallucinations Zhao et al. (2023).
Figure 1: Compared with RAG (Top) and GAG (Middle), the proposed IAG (Bottom) eschews external resources, utilizing solely LLMs to imagine a shorter explicit document and more flexible implicit adapters.
In recent years, to alleviate the issue of hallucinations in LLMs and improve performance in knowledge-intensive tasks such as QA, researchers have proposed numerous knowledge-augmented methods for LLMs, which mainly include two categories: Retrieval-Augmented-Generation (RAG) Guu et al. (2020) and Generation-Augmented-Generation (GAG) Abdallah and Jatowt (2023). RAG (Top part of Figure 1) retrieves related documents from external resources (e.g., auxiliary tools and domain documents) and then sends those retrieved documents and the question together into LLMs Izacard and Grave (2021). RAG has demonstrated formidable performance on several tasks Lewis et al. (2020). However, RAG not only requires presupposed external resources but also necessitates more computational resources and longer processing times Xu et al. (2023a). Take typical RAG method FiD Izacard and Grave (2021) as an example, the required computing resources and inference time continue to increase as the number of retrieved documents increases. While retrieving 100 documents requires processing over 12k tokens, it will result in an exceeding 100×100$\times$100 × prompt length (decades times of GPU consumption) and an over $100^{2}\times$ computational time Liu et al. (2023a).
To avoid dependence on external resources, GAG Yu et al. (2023) has been proposed and can utilize LLMs like InstructGPT Ouyang et al. (2022) to generate the relevant documents (Middle part of Figure 1). However, it requires additional financial costs (e.g., API calls) and still demands substantial computational resources and time. Additionally, both RAG and GAG utilize more explicit external resources (symbolized documents), and the quality of the obtained content significantly impacts downstream tasks Li et al. (2023); Shaier et al. (2024). For example, Gao et al. (2024) indicates that noise in the documents negatively affects the performance. Therefore, there is an urgent need to explore new knowledge-augmented methods.
@@
In fact, LLMs inherently contain rich knowledge and possess significant potential for resolving knowledge-intensive tasks Bhagavatula et al. (2020). Enhancing the performance of specific tasks can be achieved by better activating relevant knowledge or expanding memory capacity without relying on external resources. For example, simply repeating the question twice Xu et al. (2023b), just reviewing and consolidating knowledge by appending a straightforward prompt “As far as I know” Yao et al. (2023), and using visual-language models to imagine images Tang et al. (2023), all approaches enhance the performance of LLMs on downstream tasks. Inspired by this, we introduce a novel knowledge-augmented framework Imagination-Augmented-Generation (IAG) for LLMs, which simulates the human capacity to compensate for knowledge deficits solely through imagination in QA. As shown in the bottom part of Figure 1, for resolving knowledge-intensive tasks, IAG utilizes solely LLMs to imagine a shorter explicit document and more flexible implicit adapters.
@@
Within the framework of IAG, we introduce an imagine richer context method for question answering (IMcQA). To sufficiently utilize the inherent knowledge of LLMs, we design two main modules to activate the various potential knowledge modeling in LLMs and obtain a richer context. Specifically, the explicit imagination module first uses symbol distillation to obtain the compressed context and then guides LLMs in generating a short and useful dummy document. Subsequently, the implicit imagination module utilizes the proposed HyperNetwork to generate LoRA weights to activate the task-processing ability of LLMs. Unlike the LoRA Hu et al. (2021) stores task knowledge and ability in modules, the HyperNetwork learns to imagine hidden knowledge for each question.
@@
We evaluate the proposed IMcQA to various LLMs, including T5 Roberts et al. (2020a) and Llama2 Touvron et al. (2023). The experimental results across three QA datasets indicate that the proposed method yields performance gains while reducing computational expenses and time. Notably, it even outperforms baseline methods that retrieve and generate knowledge under the same document settings. In conclusion, the contributions of this paper are summarized as follows:
@@
  • We propose a novel QA method IMcQA that employs two main modules (explicit imagination and implicit imagination) to better utilize the knowledge stored in the LLMs and obtain richer context in QA.
  • Experimental results indicate that the proposed method successfully activates the relevant internal knowledge of LLMs. IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in both in-distribution performance and out-of-distribution generalizations.
@@ 2 Related Work
Knowledge Enhancement has usually been adopted to alleviate the issue of insufficient knowledge in LLMs. There are two main methods: RAG Sun et al. (2019) and GAG Abdallah and Jatowt (2023). The typical RAG method FiD Izacard and Grave (2021) retrieves the documents from an external knowledge base to answer questions. As LLMs can be considered a knowledge base, several studies Liu et al. (2022) propose to extract knowledge from LLMs (e.g., GPT-3). For example, Yu et al. (2023) generates 10 relevant documents for world knowledge according to the question. However, RAG needs the related external resources, and both RAG and GAG still need to obtain and utilize verbose explicit long contexts. Recently, there have been some methods to enhance the LLMs’ ability through simulating human imagination of visual information. But they use existing visual-language models Tang et al. (2023); Akter et al. (2024), while we prefer self-imagination to augment knowledge. Besides, they have not fully leveraged the parameterized knowledge within the models Xu et al. (2023b); Kazemnejad et al. (2023). In this paper, the proposed method for augmenting knowledge not only obviates the need for external resources but also enhances the efficiency of extracting and activating internal knowledge within LLMs.
@@
Context Compression has often been used to improve the efficiency of LLMs in processing long contexts. Recent studies Mu et al. (2023) propose that long contexts be condensed into summary vectors (soft prompts) to ensure their effective utilization by LLMs. Simultaneously, some studies Jiang et al. (2023) believe that information redundancy in lengthy texts and information entropy can be utilized to compress the contexts Li et al. (2023). Unlike them, this paper is devoted to awakening the long-context modeling ability of LLMs. By learning an Imagine Model that can generate compressed contexts, the QA model that operates on short contexts can also possess a rich contextual understanding akin to the QA model designed for processing longer contexts.
@@
Knowledge Distillation is a technique where a smaller model learns to mimic the predictions of a larger model, aiming to retain performance while reducing computational resources Hinton et al. (2015). Recent studies West et al. (2022) present symbolic knowledge distillation, a process that facilitates knowledge transfer from a teacher model via extracting training data to subsequently train a student model. In this paper, the process of obtaining compressed context during the explicit imagination resembles a form of symbolic distillation. Regarding training, our emphasis lies in distilling the long-context modeling abilities of LLMs.
@@
Hypernetworks is designed to reduce the number of parameters Ha et al. (2016), i.e., a small neural network generates parameters for another big neural network. Recent studies Phang et al. (2022); Ivison et al. (2023) have explored the enhancement of model performance in zero- and few-shot settings through meta-learning involving hypernetworks. We utilize hypernetworks to acquire implicit imagine capabilities by dynamically generating LoRA for efficiency and generalization.
@@ 3 Method
In this section, we introduce the detailed method of IMcQA to activate LLMs’ intrinsic knowledge and obtain a richer context for QA. The fundamental premise underlying this method is that QA with a richer context yields greater performance. Consequently, diverse methods are employed for questions lacking in richer contexts to activate knowledge within LLMs to replicate comparable effects to those achieved with richer contexts.
Specifically, IMcQA comprises two main modules. Explicit imagination with long context compression learns to imagine a short dummy document (§ 3.2). And implicit imagination with the HyperNetwork models hidden knowledge that learns a shared knowledge feature projection across questions (§ 3.3). The HyperNetwork is trained to generate lightweight LoRA modules, aiming to align the question and the internal knowledge. Besides, there is long context distillation in training, which learns the teacher’s rich representations to compensate for missing knowledge in imagination (§ 3.4).
@@ 3.1 Formulation
The formulation of our task follows RAG for QA Guu et al. (2020). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$.
The goal of QA is to learn a distribution function $p(a|q)$. In a closed-book setting, LLMs directly encode the given question $q$ and generate the answer $a$ Roberts et al. (2020b). However, employing a direct approach of requesting models to output answers frequently results in poor performance, primarily attributable to the omission of a substantial amount of world knowledge. Therefore, a popular approach is the open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$. Additional background details are available in A.1.
Figure 2: Overview of IMcQA method. In the inference phase, for each question, the Imagine Model imagines a short dummy document based on the question and the HyperNetwork generates specific LoRA weights. During training, there are two stages: the first stage is the pre-training of the Imagine Model, aiming at its ability to imagine a short dummy document based on the question, and the second stage is the HyperNetwork fine-tuning using long context distillation (§ 3.4) to learn a map from the question to the LoRA weights.
3.2 Explicit Imagination with Compress
To get the context $c$, we utilize LLMs to imagine a short dummy document, which can mitigate knowledge corpus error Lee et al. (2023) by considering potentially useful contexts. In the view of compression, we greatly reduce input length, minimize noise, and elevate the salience of essential tokens.
As shown in the left part of Figure 2, to help LLMs fully utilize the knowledge and imagine compressed text, we first pretrain it on our collected question-compressed document pairs. By leveraging symbolic distillation, we employ the long-context compression method LongLLMLingua Jiang et al. (2023) to condense a large corpus of retrieved documents. These compressed texts $c^{\prime}$ then serve as fine-tuning data alongside specific prompts $p_{q}$ (A.2) and question-answer pairs for the Imagine Model $G_{\theta}$ ($\theta$ represents the model’s parameter), which guides the model to think about its knowledge and imagine a short dummy document:
@@ 3.2 Explicit Imagination with Compress
where $d$ is the dummy document generated from the Imagine Model. This process enables LLMs to conceive compressed knowledge that robustly parallels the question’s knowledge requirements.
@@ 3.3 Implicit Imagination with HyperNetwork
We advance upon LoRA Hu et al. (2021) by suggesting the implementation of the HyperNetwork, which does not directly optimize the LoRA module but generates specific LoRA adapter weights using the inputs for QA (bottom part of Figure 2). This is akin to repeating the question in the prompt Xu et al. (2023b) and incorporating certain topical cues to stimulate the model’s recall of relevant questions Wang et al. (2023). However, the distinction lies in the fact that they serve as wake-up features, whereas we are generating model parameters.

### v3

@@ Abstract
Retrieval-Augmented-Generation and Gener-ation-Augmented-Generation have been proposed to enhance the knowledge required for question answering over Large Language Models (LLMs). However, the former relies on external resources, and both require incorporating explicit documents into the context, which increases execution costs and susceptibility to noise data. Recent works indicate that LLMs have modeled rich knowledge, albeit not effectively triggered or awakened. Inspired by this, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG), which simulates the human capacity to compensate for knowledge deficits while answering questions solely through imagination, thereby awakening relevant knowledge in LLMs without relying on external resources. Guided by IAG, we propose an imagine richer context method for question answering (IMcQA). IMcQA consists of two modules: explicit imagination, which generates a short dummy document by learning from long context compression, and implicit imagination, which creates flexible adapters by distilling from a teacher model with a long context. Experimental results on three datasets demonstrate that IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalization111 Our code will be available at https://anonymous.4open.science/r/IMcQA.
@@ 1 Introduction
Knowledge-intensive tasks like question answering (QA) necessitate utilizing extensive world and domain knowledge Berant et al. (2013); Joshi et al. (2017); Kwiatkowski et al. (2019). Nowadays, Large Language Models (LLMs) have displayed notable competencies in almost every task and industry Liu et al. (2023b). However, LLMs lack the sufficient capability to independently handle knowledge-intensive tasks Frisoni et al. (2024) and usually generate hallucinations Zhao et al. (2023).
Figure 1: Compared with RAG and GAG , the proposed IAG eschews external resources, generates a shorter context (explicitly imagination) and creates flexible adapters (implicitly imagination) for each question.
In recent years, to mitigate hallucinations in LLMs and improve performance in knowledge-intensive tasks such as QA, researchers have proposed numerous knowledge-augmented methods for LLMs. These methods primarily fall into two categories: Retrieval-Augmented-Generation (RAG) Guu et al. (2020) and Generation-Augmented-Generation (GAG) Kim et al. (2024). RAG (Top part of Figure 1) retrieves related documents from external resources (e.g., Wikipedia) and then inputs both the retrieved documents and the question into LLMs Izacard and Grave (2021). GAG (Middle part of Figure 1) leverages external LLMs such as ChatGPT Ouyang et al. (2022) to generate relevant documents Yu et al. (2023).
However, these methods have the following disadvantages [^2] A more intuitive comparison can be seen in A.1.: 1) Dependence on external resources, RAG relies on external domain knowledge resources Ke et al. (2024), while GAG depends on a more powerful external LLM as a knowledge generator. This dependency hinders their broader application. 2) Increased execution costs, The required computing resources and inference time continue to increase as the number of documents increases Xu et al. (2023a). For example, the typical RAG method FiD Izacard and Grave (2021) needs to handle over 12K tokens for retrieving 100 documents, leading to more than a $100$-fold increase in prompt length and over $100^{2}$-fold increase in inference time Liu et al. (2023a). The GAG method Yu et al. (2023) requires additional financial costs (e.g., API calls). 3) Susceptibility to noise data, both methods are prone to the inclusion of irrelevant or noisy data Li et al. (2023); Shaier et al. (2024). For example, Gao et al. (2024) indicates that noise in the documents negatively affects the performance. Therefore, there is an urgent need to explore new knowledge-augmented methods.
@@
In fact, LLMs inherently contain rich knowledge and possess significant potential for resolving knowledge-intensive tasks Bhagavatula et al. (2020). Enhancing the performance of specific tasks can be achieved by better activating and awakening relevant knowledge or expanding memory capacity without relying on external resources. For example, simply repeating the question twice Xu et al. (2023b), just reviewing and consolidating knowledge by appending a straightforward prompt “As far as I know” Yao et al. (2023), and using visual-language models to imagine images Tang et al. (2023), all these approaches enhance the performance of LLMs on downstream tasks.
Therefore, inspired by the above phenomena and aiming to address the challenges faced by RAG and GAG, we propose a novel knowledge-augmented framework, Imagination-Augmented-Generation (IAG). IAG simulates the human capacity to compensate for knowledge deficits solely through imagination in QA. For resolving knowledge-intensive tasks, IAG utilizes solely LLMs to generate a shorter context and create a more flexible adapter (Bottom part of Figure 1). Through a knowledge distillation framework, IAG achieves knowledge enhancement through imagination. The fundamental idea is to permit student models, which are deficient in rich contextual information (e.g., without abundant retrieved documents), to imitate teacher models that are endowed with such information.
@@
Within the framework of IAG, we introduce an imagine richer context method for question answering (IMcQA). To sufficiently utilize the inherent knowledge of LLMs, we design two main modules to activate the various potential knowledge modeling in LLMs and obtain a richer context. Specifically, the explicit imagination module first uses symbol distillation to obtain the compressed context and then guides the context generator to generate a short dummy document. Subsequently, within the knowledge distillation framework, the implicit imagination module utilizes the hypernetwork to generate adapters to awaken the question-aware ability of LLMs. Unlike the LoRA Hu et al. (2021) stores task knowledge and ability in modules, the hypernetwork learns to generate a question-aware knowledge module.
@@
We evaluate the proposed IMcQA on various LLMs, including T5 Roberts et al. (2020a) and Llama2 Touvron et al. (2023). The experimental results across three QA datasets indicate that the proposed method yields performance gains while reducing computational expenses and time. Notably, it even outperforms baseline methods that retrieve and generate knowledge under the same document settings. In conclusion, the contributions of this paper are summarized as follows:
@@
  • We propose a novel QA method IMcQA that employs two main modules (explicit imagination and implicit imagination) to utilize better the knowledge stored in the LLMs and awaken a richer context in QA.
  • Experimental results indicate that the proposed method successfully activates the relevant internal knowledge of LLMs. IMcQA exhibits significant advantages in both open-domain and closed-book settings, as well as in out-of-distribution generalizations.
@@ 2 Related Work
Knowledge Enhancement has usually been adopted to alleviate the issue of insufficient knowledge in LLMs. There are two main methods: RAG Sun et al. (2019); Wang et al. (2024) and GAG Abdallah and Jatowt (2023). The typical RAG method FiD Izacard and Grave (2021) retrieves documents from Wikipedia to answer questions. LLMs serving as a knowledge base have been the focus of numerous studies that advocate the extraction of knowledge from such models (e.g., GPT-3). For instance, Yu et al. (2023) generates 10 documents for each question. However, RAG requires external resources, and both RAG and GAG need verbose long contexts. Recently, methods have been developed to enhance LLMs’ abilities by simulating human imagination of visual information using existing visual-language models Tang et al. (2023); Akter et al. (2024). We prefer self-imagination to augment knowledge and aim to leverage the parameterized knowledge within the models Xu et al. (2023b); Kazemnejad et al. (2023). Our proposed method not only eliminates the need for external resources but also improves the efficiency of activating internal knowledge within LLMs.
@@
Context Compression has often been used to improve the efficiency of LLMs in processing long contexts. Recent studies Mu et al. (2023) propose that long contexts be condensed into summary vectors (soft prompts) to ensure their effective utilization by LLMs. Simultaneously, some studies Jiang et al. (2023); Pan et al. (2024) suggest utilizing information redundancy and entropy in lengthy texts to compress contexts Li et al. (2023). Unlike these approaches, this paper aims to enhance the long-context modeling ability of LLMs. By developing a context generator that creates compressed contexts, the QA model operating on short contexts can achieve a rich contextual understanding similar to models designed for longer contexts.
@@
Knowledge Distillation is a technique where a smaller model learns to mimic the predictions of a larger model, aiming to retain performance while reducing computational resources Hinton et al. (2015). Recent studies West et al. (2022) present symbolic knowledge distillation, a process that facilitates knowledge transfer from a teacher model via extracting training data to subsequently train a student model Wang et al. (2023b); Ranaldi and Freitas (2024). In this paper, the process of obtaining compressed context during the explicit imagination resembles a form of symbolic distillation. Regarding training, our emphasis lies in distilling the long-context modeling abilities of LLMs.
@@
Hypernetworks is designed to reduce the number of parameters Ha et al. (2016), i.e., a small neural network generates parameters for another big neural network. Recent studies Phang et al. (2022); Ivison et al. (2023) have explored the enhancement of model performance in zero/few-shot settings through meta-learning involving hypernetworks. We utilize hypernetworks to acquire implicit imagine capabilities by dynamically generating LoRA for efficiency and generalization.
@@ 3 Method
In this section, we introduce the detailed method of IMcQA to activate LLMs’ intrinsic knowledge and obtain a richer context for QA. The fundamental premise underlying this method is that QA with a richer context yields greater performance (e.g., RAG with retrieved documents). Consequently, diverse methods are employed for questions lacking in richer contexts to activate knowledge within LLMs to replicate comparable effects to those achieved with richer contexts.
Specifically, IMcQA comprises two main modules. Explicit imagination with long context compression learns to imagine a short dummy document (§ 3.2). And implicit imagination with the hypernetwork models’ hidden knowledge that learns a shared knowledge feature projection across questions (§ 3.3). The hypernetwork is trained to generate lightweight LoRA modules, aiming to align the question and the internal knowledge. Besides, there is long context distillation in training, which learns the teacher’s rich representations to compensate for missing knowledge in imagination (§ 3.4).
@@ 3.1 Formulation
The formulation of our task follows RAG for QA Guu et al. (2020). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$. Additional background details are available in B.1.
Figure 2: Overview of IMcQA method. In the inference phase, for each question, the explicit imagination generates a short dummy document and the implicit imagination with hypernetwork creates a specific LoRA module. During training, there are two stages: the first stage is the pre-training of the context generator (§ 3.2), aiming at its ability to imagine a short dummy document based on the question, and the second stage is the hypernetwork fine-tuning (§ 3.3) using long context distillation (§ 3.4) to obtain a question-specific LoRA module.
3.2 Explicit Imagination with Context Generator
To obtain a context $c$ for each question, we utilize LLMs to generate a short dummy document, which can mitigate knowledge corpus error Lee et al. (2023) by considering potentially useful contexts. We employ the "teacher-student" framework. During the training phase, the teacher model, equipped with rich contextual information, generates short contexts through context compression. The student model, in turn, learns to automatically generate these short contexts by leveraging its own LLM capabilities.
As shown in the left part of Figure 2, to help LLMs fully utilize the knowledge and generate compressed text, we first pretrain the context generator on our collected question-compressed document pairs. By leveraging symbolic distillation, we employ the long context compression method LongLLMLingua Jiang et al. (2023) to condense a large corpus of retrieved documents. These compressed texts $c^{\prime}$ then serve as fine-tuning data alongside specific prompts $p_{q}$ (B.2) and question-answer pairs for the context generator [^3] We discuss the role of context generator in the A.2. $G_{\theta}$ ($\theta$ represents the model’s parameter), which guides the model to think about its knowledge and generate a short dummy document:
@@ 3.2 Explicit Imagination with Context Generator
where $d$ is the dummy document generated from the context generator. This process enables LLMs to conceive compressed knowledge that robustly parallels the question’s knowledge requirements.
@@ 3.3 Implicit Imagination with Hypernetwork
Generally speaking, richer context can help LLM better answer questions. That is, the representation of questions and the internal state of LLM when utilizing rich context are the better states. Therefore, in the absence of context, we should focus on building models to awaken LLM to achieve this better state and as a better QA model.
We utilize the hypernetwork [^4] We conduct a detailed analysis of the reasons behind the hypernetwork in the A.3. to learn to generate a better QA model, which does not directly optimize the LoRA Hu et al. (2021) module but generates specific LoRA adapter weights using the inputs for QA (bottom part of Figure 2). This is akin to repeating the question in the prompt Xu et al. (2023b) and incorporating certain topical cues to stimulate the model’s recall of relevant questions Wang et al. (2023c). However, the distinction lies in the fact that they serve as wake-up features, whereas we are generating model parameters.
@@ 3.4 Training with Long Context Distillation
Models # Docs NQ TriviaQA WebQ
Large (800M) XL (3B) Large (800M) XL (3B) Large (800M) XL (3B)
# Closed-book Setting
T5 Roberts et al. (2020a) 0 28.5 28.30 28.7 33.92 30.6 34.43
LoRA Hu et al. (2021) 0 17.70 23.15 23.87 32.16 29.13 35.24
IMcQA (Ours) 0 29.32 29.59 30.11 35.71 32.68 37.40
# Retrieval Augmented Setting (compared with RAG)
DPR Karpukhin et al. (2020) (110M) 100 41.5 - 56.8 - 41.1 -
RAG Lewis et al. (2020) 10 44.5 - 56.1 - 45.2 -
FiD Izacard and Grave (2021) 10 46.7 50.1 61.9 66.3 48.1 50.8
FiD Izacard and Grave (2021) 100 51.4 55.2 67.6 72.9 50.5 52.9
EAR Chuang et al. (2023) 10 39.6 42.3 60.0 64.6 - -
RFiD Wang et al. (2023a) 10 48.3 50.5 63.4 67.8 - -
FILCO Wang et al. (2023d) 1 - 44.7 - 59.0 - -
IMcQA (Ours) 10 49.9 50.9 69.7 70.3 51.5 52.8
IMcQA (Ours) 30 53.1 - 70.5 - 52.0 -
# Generation Augmented Setting (compared with GAG)
GENREAD (sampling) Yu et al. (2023) 10 40.3 42.6 67.8 69.6 51.5 52.6
GENREAD (clustering) Yu et al. (2023) 10 43.5 45.6 70.2 71.6 53.5 54.4
IMcQA (Ours) 10 48.8 49.2 70.9 72.2 54.5 55.6
# Imagination Augmented Setting
LoRA Hu et al. (2021) 1 40.1 44.2 62.8 66.9 43.7 48.2
IMcQA (Ours) 1 42.3 46.5 65.5 68.4 45.3 50.5
Table 1: QA performances of different methods with different settings. The first part (closed-book setting) indicates that only utilize questions; The latter three parts utilize explicit documents. The best results are in bold, while the second-best are underlined. * means that those results are from existing papers, denotes that the documents were generated ( indicates that the number of documents is reduced due to insufficient memory for distillation).